Einstein Inc. is evaluating different equipment. Machine A costs $55,000 per year has a five year life, and costs $15,000 pear year to operate. The machine will be depreciated using straight line and the relevant discount rate is 10%. The machine will have a salvage value of $8,500 at the end of the projects life. The firm has a tax rate of 21%.

a.) What is the operating cash flow in year 1? (Enter a negative value)

b.) What is the NPV of the project? (Enter a negative value and round to 2 decimals)

c.) What is the EAC for the project? (Enter a positive value and round to 2 decimals)
Answer #1

1.
=(-15000-55000/5)*(1-21%)+55000/5=-9540

2.
=-55000+((-15000-55000/5)*(1-21%)+55000/5)/10%*(1-1/1.1^5)+8500*(1-21%)/1.1^5=-86994.6190958144

3.
=-(-55000+((-15000-55000/5)*(1-21%)+55000/5)/10%*(1-1/1.1^5)+8500*(1-21%)/1.1^5)*10%/(1-1/1.1^5)=22948.9613601743